in the last five seconds of the championship match between teams A and B, the scores were tied at 75 when a player was fouled and was awarded two free throws. On the first attempt, he released the ball with a velocity of 6.0 m/s at 45 degerees above the horizontal from a height of 2.0 m from the floor.
a.)Is this a miss or a basket?
b.) For the second throw, the player shoots again at 45 degrees from the same point of release. What must the initial speed of the ball be so that it will go through thr basket?
1
Expert's answer
2019-07-22T11:33:36-0400

The answers are:


a) miss

b) more than 6.3 m/s, because


the ball from the point of throw should rise by 1 m, what means

"v^{2} = \frac{2gh}{sin^{2}45}"

correspondingly minimum speed

"v = 2(gh)^{1\/2} = 6.3 (m\/s)"